OpenStudy (anonymous):

A student found the slope of a line that passes through the points (3, 15) and (5, 9) to be -1/3. What mistake did he make? A. He used the negative reciprocal of the slope. B. He added a negative sign. C. He used the reciprocal of the slope. D. He did not make a mistake. That is the correct slope.

OpenStudy (anonymous):

@MrNood @Ms-Brains @nincompoop

OpenStudy (anonymous):

@retirEEd hii

OpenStudy (retireed):

Can you calculate the correct slope?

OpenStudy (anonymous):

yes?

OpenStudy (retireed):

What is it?

OpenStudy (anonymous):

-1/3?

OpenStudy (retireed):

No, that was the incorrect slope use .... slope = (y2-y1) / (x2-x1)

OpenStudy (anonymous):

-3?

OpenStudy (retireed):

That's correct. So what mistake did the student make?

OpenStudy (anonymous):

C?

OpenStudy (retireed):

That's what I think.

OpenStudy (anonymous):

medal (:

OpenStudy (anonymous):

can u help with a couple more?

OpenStudy (retireed):

If you show me your work I will continue to help.

OpenStudy (anonymous):

ok (:

OpenStudy (anonymous):

What is the point-slope equation of the line with slope -12 that goes through the point (5, 3)? A. y - 3 = -12(x - 5) B. y + 3 = -12(x + 5) C. y + 3 = 12(x + 5) D. y - 3 = 12(x - 5)

OpenStudy (anonymous):

A?

OpenStudy (retireed):

Why? I am not very familiar with this form. I like slope intercept form better.

OpenStudy (retireed):

i see it is going to be either A or B since I see the -12x.

OpenStudy (retireed):

I'm trying to graph these on my TI-84

OpenStudy (anonymous):

yea i GF said it was A

OpenStudy (anonymous):

so i think that is the correct one

OpenStudy (retireed):

I got A also.

OpenStudy (retireed):

Are you sure you need help? You are better at these than me.

OpenStudy (anonymous):

lol i like the have a second opinion on things